the manager of a bulk foods establishment sells a trail mix for $9 per pound and premium cashews for $14 per pound. The manager wishes to make a 225-pound trail mix-cashew mixture that will sell for $11 per pound. How many pounds of each should be used?
Answer by scott8148(6628)

c + t = 225

9t + 14c = 11(225)

solve the system using substitution or elimination